OpenStudy (momo):

The slant height of a regular pentagonal pyramid is 8 cm, and the length of each side of the base is 4 cm. Find the lateral area.

OpenStudy (anonymous):

LA = (1/2)(perimeter of base)(slant height) LA = (1/2)(2*(4+4))(8) LA = (1/2)(16)(8) LA = 64 cm

OpenStudy (momo):

could that be reduced? because 64 isn't an option for me...

OpenStudy (momo):

here are my options, 32, 60, 80, 160

OpenStudy (anonymous):

Sorry i made a stupid mistake one second

OpenStudy (momo):

oh it's okay :)

OpenStudy (anonymous):

LA = (1/2)(perimeter)(base length) LA = (1/2)((4+4) * 2)(4) LA = 32

OpenStudy (momo):

thanks!!